OpenStudy (anonymous):

what is the derivative of: A = (3pi)(9+h)(sqrt(18-2h))

OpenStudy (anonymous):

\[A = (3\Pi)(9+h)(\sqrt{18-2h})\]

OpenStudy (anonymous):

@sami-21

OpenStudy (anonymous):



OpenStudy (anonymous):

use product rule where 3 pi is constant \[\Large \frac{d}{dh}(f*g)=\frac{d}{dh}(f)*g+\frac{d}{dh}(g)*f\]

OpenStudy (anonymous):

here f=(9+h) g=sqrt(18-2h) does that help ?

OpenStudy (anonymous):

o...k...

OpenStudy (anonymous):

is the derivative of 9+h=1?

OpenStudy (anonymous):

exactly !

OpenStudy (anonymous):

okay..! thanks a lot! :)

OpenStudy (anonymous):

you are welcome :)